A binomial experiment with probability of success p=0.73 and n=7 trials is conducted. What is the probability that the experiment results in 6 or more successes?
Answer by Theo(13342)

p(x) = p^x * q^(n-x) * c(n,x)

the sum of all probabilities needs to be 1.

in your problem:

p = .73
q = 1 - .73 = .27

p(6) = .73^6 * .27^1 * c(7,6) = .2860216877
p(7) = .73^7 * .27^0 * c(7,7) = .1104739852

your solution is the sum of these = .3964956729.

c(n,x) is equal to n! / (x! * (n-x)!)

for example:
c(7,6) is equal to 7! / (6! * 1!) = (7 * 6!) / (6! * 1!) = 7

here's all the probabilities that i calculated using excel.
the total probability is 1 as it should be.